# LoRA is All You Need for Safety Alignment of Reasoning LLMs

### v3

@@ 4.3 Datasets and Models
Models. We conduct experiments on two widely used open-weight reasoning-capable models: DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-14B. Safety evaluation is performed using Llama-Guard-3-8B, which was found to be the most accurate evaluator in (Jiang et al., 2025).
Safety fine-tuning dataset. We use the DirectRefusal dataset, adapted from (Rosati et al., 2024) by (Huang et al., 2025), which provides harmful requests paired with refusal-style answers.
Safety evaluation dataset. We adopt the StrongREJECT dataset (Souly et al., 2024), which consists of 310 policy-violating queries designed to test whether the model behaves safely.
@@
Training Setup. Full-model fine-tuning is performed for 5 epochs, while LoRA fine-tuning is run for 10 epochs. We save and evaluate checkpoints at every epoch. Unless otherwise stated, LoRA is applied only to the MLP layers with rank $r=1$. In Section 5, we investigate the effect of varying $r$ and applying LoRA to different modules (e.g., MLP vs. attention) and layers.
Additional experimental details are deferred to Appendix A.
@@ 4.4 LoRA is All You Need for Safety Alignment of Reasoning LLMS
(a) DeepSeek-R1-Distill-Qwen-7B
(b) DeepSeek-R1-Distill-Qwen-14B
Figure 3: LoRA bypasses the “Safety Tax”, achieving safety comparable to that of the full-model fine-tuned model and reasoning performance comparable to the original reasoning model. We plot reasoning performance—measured by Pass@1—against safety scores for different models. For the fine-tuned models, we report results for checkpoints at all epochs. Results on the base versions of HumanEval and MBPP are provided in Figure 11 in the Appendix, where the same patterns hold, but with higher accuracy.
Figure 3 compares the safety and reasoning capabilities at different checkpoints (i.e., epochs), during full-model and LoRA safety alignment fine-tuning. We observe that the base model before safety fine-tuning exhibits high accuracy but low safety. On the other hand, the fully fine-tuned models achieve good safety at the cost of reduced accuracy. In contrast, the LoRA fine-tuned models maintain strong performance in both safety and reasoning (as evidenced by the red points in the upper-right corner of the plot).
For the 7B models, the best LoRA checkpoints outperform the base model on all tasks. In terms of safety, they fall slightly short of the fully fine-tuned models, with an average drop of about 0.03 in safety score. For the 14B models, there is a consistent but small decrease in reasoning accuracy compared to the base model, while safety performance remains comparable to that of the fully fine-tuned model.
@@ 5 How Much LoRA is Enough?
In this section, we ablate the LoRA configuration to identify the key elements that matter most. We examine three factors: (1) the rank $r$, (2) the modules to which LoRA is applied, and (3) the layers to which LoRA is applied. Our goal is to determine the minimal setup needed for LoRA—i.e., the smallest update sufficient to achieve both strong reasoning and safety.
To compare different LoRA configurations, we visualize performance in the reasoning–safety plane. For each configuration, we train for 10 epochs, which yields multiple checkpoints with varying reasoning and safety scores. We then evaluate configurations by examining the Pareto frontier of these checkpoints, which allows us to compare tradeoffs effectively.
@@ 5.1 Rank: $r=1$ is Sufficient
We explore the effect of the rank $r$ in LoRA. Specifically, we run experiments with the 14B model by applying LoRA to the MLP layers while varying $r$. Results at the final checkpoint are shown in Figures 4(a) and 4(b). We also include full-model fine-tuning as the rightmost point, since it allows for full-rank updates. Reasoning performance generally declines as $r$ increases, although the drop is minor between $r=1$ and $r=8$. For safety scores, we observe a notable decrease when $r$ increases from 4 to 64, whereas full-model fine-tuning yields a better safety score than $r=64$. We suspect this may be due to optimization difficulties at intermediate high ranks, while very low-rank (under-parameterized) or full-rank (over-parameterized) setups may benefit from easier optimization dynamics, leading to better safety outcomes. Most importantly, we find that $r=1$ is sufficient to achieve the best performance in both reasoning and safety. Figure 4(c) further confirms that $r=1$ also achieves the best (on AIME) or nearly best (on GPQA) tradeoff when varying the number of training epochs. This indicates that we can achieve strong performance on both reasoning and safety at the lowest fine-tuning cost.
Additionally, the fact that $r=1$ is sufficient reflects the inherently low-rank nature of the safety alignment task itself. This connects to prior work showing that safety can be mediated by a single direction within the model’s internal representations, often referred to as a steering vector (Panickssery et al., 2023) or refusal features (Arditi et al., 2024; Yu et al., 2024). This perspective may explain why a rank-1 update is enough to achieve safety.
(a) Reasoning performance vs. $r$
(b) Safety v.s. $r$
(c) Reasoning-safety tradeoff
Figure 4: In (a) and (b), we show reasoning and safety performance at different LoRA ranks $r$ for the 14B model, respectively. Full-model fine-tuning is included as the rightmost point for reference. Reasoning performance decreases as $r$ increases, while safety first decreases and then increases. Overall, very low ranks are recommended, and $r=1$ already achieves the best performance in both metrics. In (c), we visualize the Pareto frontiers of the reasoning–safety tradeoff when the training epoch is varied and observe that $r=1$ is sufficient to achieve an excellent tradeoff, outperforming other ranks, especially on AIME.
@@ 5.2 Modules: Up-Projection Matters Most in MLPs
The LoRA adapter is usually applied to attention layers and/or MLP modules, e.g., by setting target_modules when using the PEFT package. Here, we explore the effect of applying LoRA to different modules. We first compare applying it to both attention and MLP layers (QKVO & MLP in the figure) versus only applying it to MLP layers. Figure 5 shows the comparison for the 14B models with $r=1$. We observe that applying LoRA only to the MLP layers yields a similar Pareto frontier compared to applying it to both modules.
Next, we perform an ablation over the modules within the MLP. Specifically, in the Qwen architecture, the MLP layers use the popular SwiGLU Chowdhery et al. (2022) structure that contains a gate projection, an up projection, and a down projection. We apply LoRA to only one of these projection layers at a time and evaluate the results. We use $r=1$. Figure 6 shows that across different tasks, applying LoRA only to the up projection achieves strong results, with the Pareto frontier often on par with using the full MLP and even outperforming it on the two coding benchmarks. In contrast, applying LoRA only to the down projection yields noticeably worse performance. These findings suggest that different projections within the MLP contribute differently to the reasoning–safety tradeoff, and that the up projection is particularly important and sufficient by itself in our setup.
@@
Discussion. We mainly focus our ablations on the MLP, as it is primarily responsible for feature transformations and thus well-suited for safety alignment. However, we believe it is also worthwhile to study ablations over the attention layers, which we leave for future work. It is also valuable to further investigate why the up projection alone is the best choice. Additionally, extending these experiments to other LLM architectures would be an important direction for future work.
@@ 5.3 Layers: Middle Layers Matter Most
We ablate over layers in the model. The 14B Qwen2.5 architecture has 48 layers in total, and we apply LoRA to only 16 of them. We consider three configurations: (1) layers 5–20, denoted as “Early Layers”, (2) layers 17–32, denoted as “Middle Layers”, and (3) layers 25–40, denoted as “Late Layers”. In all cases, we apply LoRA only to the up projection layers with $r=1$. The results are shown in Figure 7. Across tasks, we observe that applying LoRA to the middle layers yields the best tradeoff, achieving performance on par with using all layers on AIME and GPQA, and only slightly behind using all layers on HumanEval+ and MBPP+. In contrast, applying LoRA to either the early or late layers results in a noticeably worse tradeoff. This shows that the middle layers are most important for balancing reasoning and safety. Interestingly, this again connects to prior findings on steering vectors (Panickssery et al., 2023) and refusal features (Arditi et al., 2024; Yu et al., 2024), which suggest that directions in intermediate representations responsible for safety behavior are most prominent in the middle layers.
6 Exploring the Structure of LoRA Weights
(a) $\frac{\|{\bm{W}}_{I}^{\top}\Delta{\bm{W}}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|}$
(b) $\frac{\|{\bm{U}}_{16}{\bm{U}}_{16}^{\top}\Delta{\bm{W}}\|}{\|\Delta{\bm{W}}\|}$
(c) $\frac{\|{\bm{W}}_{I}\Delta{\bm{W}}^{\top}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|}$
Figure 8: LoRA updates exhibit smaller overlap with the original weights compared to the full-model fine-tuning updates. Although the reduction in overlap is sometimes small, it can be observed across most layers for all four metrics, which cover both the column (a)(b) and row spaces (c)(d). The 14B models are used here.
6.1 LoRA Updates Have Less Alignment with Initial Weights
We compare the full-rank fine-tuned model with the LoRA fine-tuned model in which both the attention and MLP modules are updated with $r=4$, making the two settings more comparable since updates occur in all major modules. We compute the alignment metrics for different module types across layers, average them over layers, and report the results for the 14B models in Figure 8. We observe that LoRA achieves smaller overlap in most modules across the metrics, with a few exceptions. This suggests that, for most weights, ${\bm{W}}_{I}$ and $\Delta{\bm{W}}$ are more orthogonal—both in the column and row spaces—for LoRA than for full-model fine-tuning. In other words, under LoRA fine-tuning, the safety-oriented updates read from and write into subspaces that are more separate from those used by the original reasoning-related weights—more so than in the full-model fine-tuned version. Although the reduction in alignment values is sometimes small, it may still indicate that LoRA updates interfere less with the reasoning-related components of the model, potentially explaining the better preservation of reasoning performance. A more in-depth investigation is needed to fully understand the underlying mechanisms and to develop more precise metrics for capturing this effect—an important direction for future work.
(a) 14B models on AIME and GPQA.
(b) 14B models on HumanEval+ and MBPP+.
6.2 Exploring Methods That Further Reduce Alignment
Given the observations in Section 6.1, we ask whether further reducing the overlap between $\Delta{\bm{W}}$ and ${\bm{W}}_{I}$ could lead to even better reasoning performance without compromising safety. This question is particularly relevant because, while the results achieved by LoRA in Section 4.4 are strong, they are not perfect—a small performance gap remains compared to the original reasoning model, especially for the 14B model on AIME, HumanEval+, and MBPP+.
We experiment with two approaches: (1) Regularization during LoRA training: adding a penalty to discourage overlap between ${\bm{W}}_{I}$ and $\Delta{\bm{W}}$, targeting either the column space (reg-col) or both the column and row spaces (reg-both). For efficiency, we approximate ${\bm{W}}_{I}$ via a truncated SVD. We tried different values of $\beta$ but observed negligible differences, so we fix $\beta=1$. (2) Post-hoc Orthogonalization: post-processing $\Delta{\bm{W}}$ before applying Equation 1, projecting it onto the orthogonal complement of the top-$k$ singular vectors of ${\bm{W}}_{I}$. We test both column-only (OrthoMerge-col) and column+row (OrthoMerge-both) variants, with scaling $\lambda$ to offset safety loss. Further details are deferred to Appendix B.
Figure 9(a) shows the results on AIME and GPQA for the 14B models with $r=4$. For each method that yields multiple checkpoints (e.g., different epochs for vanilla and regularization fine-tuning, or different hyperparameters for OrthoMerge), we visualize the Pareto-frontier points. We observe that methods addressing both the row and column spaces (“both”) tend to yield better results than those that only operate on the column space (“col”). Among them, OrthoMerge-both appears most promising, with more points concentrated in the upper-right corner and its best point strictly dominating vanilla LoRA. Therefore, we additionally evaluate OrthoMerge-both on the coding benchmarks in Figure 9(b), where it slightly underperforms on HumanEval+ but slightly outperforms vanilla on MBPP+ in terms of the tradeoff.

### v4

@@ 4.3 Datasets and Models
Models. We consider three widely used open-weight reasoning models: DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Qwen-14B, and DeepSeek-R1-Distill-Llama-8B. These models span three sizes and two architectures (Qwen and Llama). Safety evaluation is performed using Llama-Guard-3-8B, which Jiang et al. (2025) found to be the most accurate safety evaluator.
Safety finetuning dataset. We use the DirectRefusal dataset, adapted from (Rosati et al., 2024) by (Huang et al., 2025), containing harmful requests with refusal answers.
Safety evaluation datasets. We consider two datasets for safety evaluation. The first is StrongREJECT (Souly et al., 2024), which consists of 310 policy-violating queries designed to test whether a model behaves safely. The second is BeaverTails (Ji et al., 2023), which contains harmful data spanning 14 harm categories. Together, these datasets provide a comprehensive evaluation of model safety. All results in the main paper are reported on StrongREJECT, and we defer results on BeaverTails to Appendix B.1. The patterns observed in our experiments are consistent across both datasets. We do not consider jailbreaking benchmarks, as they are out of scope. Jailbreaking typically requires defenses beyond safety-alignment finetuning, such as input preprocessing (Jain et al., 2023; Wang et al., 2024b; Xie et al., 2023), external detectors/classifiers (Alon and Kamfonas, 2023; Sharma et al., 2025), or inference-time interventions (Xu et al., 2024; Zou et al., 2023). Accordingly, we only focus on the capabilities that safety alignment is intended to address.
@@
Training Setup. Full-model finetuning is performed for 5 epochs and LoRA is run for 10 epochs. We save and evaluate checkpoints at every epoch. Unless otherwise stated, LoRA is applied only to the MLP layers with rank r=1r\!=\!1. In Section 5, we investigate the effect of varying $r$ and applying LoRA to different modules (e.g., MLP vs. attention) and layers. Additional details are deferred to Appendix A.
@@ 4.4 LoRA is All You Need for Safety Alignment of Reasoning LLMS
(a) R1-Distill-Qwen-7B
(b) R1-Distill-Qwen-14B
(c) R1-Distill-Llama-8B
Figure 3: Across different model sizes and architectures, LoRA bypasses the “Safety Tax”, achieving safety comparable to full-model finetuning while preserving reasoning performance close to that of the original reasoning model. We plot reasoning performance—measured by Pass@1—against safety scores for different models. For finetuned models, we report results for checkpoints from all epochs. Results on the base versions of HumanEval and MBPP are shown in Figure 13; the same pattern holds, but with higher accuracy. All safety scores here are measured on StrongREJECT; results on BeaverTails are deferred to Figure 11, where we observe the same pattern.
Figure 3 compares safety and reasoning performance across checkpoints (i.e., epochs) during full-model and LoRA-based safety alignment. We observe that the base model (before safety finetuning) achieves high accuracy but low safety. In contrast, full-model finetuning substantially improves safety at the cost of reduced reasoning accuracy. LoRA finetuning largely avoids this trade-off, maintaining strong performance on both safety and reasoning (as evidenced by the red points in the upper-right region of the plot). This pattern holds consistently across three model sizes, two architectures, two safety benchmarks, and four reasoning benchmarks spanning mathematics, science, and code generation. In some settings—e.g., the 14B model on the two coding tasks or the 8B model on AIME—we observe a slight drop in reasoning performance relative to the base model. In other cases, LoRA finetuning even improves reasoning performance (e.g., the 7B model across tasks). Overall, the trend is robust and highlights the effectiveness of LoRA compared to full-model finetuning.
@@ 5 How Much LoRA is Enough?
In this section, we ablate the LoRA configuration to identify the key elements that matter most. We examine three factors: (1) the rank $r$, (2) the modules and (3) the layers to which LoRA is applied. Our goal is to determine the minimal setup needed for LoRA—i.e., the smallest update sufficient to achieve both strong reasoning and safety.
To compare LoRA configurations, we visualize performance in the reasoning–safety plane. For each configuration, we train for 10 epochs, producing multiple checkpoints with varying reasoning and safety scores. We evaluate each configuration via the Pareto frontier of its checkpoints, enabling a clean visualization and effective comparison.
@@ 5.1 Rank: $r=1$ is Sufficient
Figure 4: We visualize the Pareto frontiers of the reasoning–safety tradeoff when the training epoch is varied and observe that $r=1$ is sufficient to achieve an excellent tradeoff, outperforming other ranks, especially on AIME.
We explore the effect of the LoRA rank $r$ by applying LoRA to the MLP layers of the 14B model while varying $r$. We find that $r=1$ is sufficient to achieve the best performance in both reasoning and safety. As shown in Figure 4, $r=1$ attains the best (on AIME) or near-best (on GPQA) trade-off across training epochs. This result suggests that the safety alignment task is inherently low-rank. Indeed, prior work has shown that safety behavior can often be mediated by a single direction in a model’s internal representations, commonly referred to as a steering vector (Panickssery et al., 2023) or refusal features (Arditi et al., 2024; Yu et al., 2024), which may explain why a rank-1 update suffices for effective safety alignment.
@@ 5.2 Modules: Up-Projection Matters Most in MLPs
The LoRA adapter is usually applied to attention layers and/or MLP modules, e.g., by setting target_modules when using the PEFT package. Here, we explore the effect of applying LoRA to different modules. We first compare applying it to both attention and MLP layers (QKVO & MLP in the figure) versus only applying it to MLP layers. Figure 5 shows the comparison for the 14B models with $r=1$. We observe that applying LoRA only to the MLP layers yields a similar Pareto frontier compared to applying it to both.
Next, we perform an ablation over the modules within the MLP. Specifically, in the Qwen architecture, the MLP layers use the popular SwiGLU (Chowdhery et al., 2022) structure that contains a gate projection, an up projection, and a down projection. We apply LoRA to only one of these projection layers at a time and evaluate the results. We use $r=1$. Figure 6 shows that across different tasks, applying LoRA only to the up projection achieves strong results, with the Pareto frontier often on par with using the full MLP and even outperforming it on the two coding benchmarks. In contrast, applying LoRA only to the down projection yields noticeably worse performance. These findings suggest that different projections within the MLP contribute differently to the reasoning–safety tradeoff, and that the up projection is particularly important and sufficient by itself in our setup.
Discussion. We mainly focus our ablations on the MLP, as it is primarily responsible for feature transformations and thus well-suited for safety alignment. We believe it is also worthwhile to study ablations over the attention layers, which we leave for future work. It is also valuable to further investigate why the up projection alone is the best choice. Additionally, extending these experiments to other LLM architectures would be an important direction for future work.
@@ 5.3 Layers: Middle Layers Matter Most
We ablate over layers in the model. The 14B Qwen2.5 architecture has 48 layers in total, and we apply LoRA to only 16 of them. We consider three configurations: (1) layers 5–20, denoted as “Early Layers”, (2) layers 17–32, “Middle Layers”, and (3) layers 25–40, “Late Layers”. In all cases, we apply LoRA only to the up projection layers with $r=1$. The results are shown in Figure 7. Across tasks, we observe that applying LoRA to the middle layers yields the best tradeoff, achieving performance on par with using all layers on AIME and GPQA, and only slightly behind using all layers on HumanEval+ and MBPP+. In contrast, applying LoRA to either the early or late layers results in a noticeably worse tradeoff. This shows that the middle layers are most important for balancing reasoning and safety. Interestingly, this again connects to prior findings on steering vectors (Panickssery et al., 2023) and refusal features (Arditi et al., 2024; Yu et al., 2024), which suggest that directions in intermediate representations responsible for safety behavior are most prominent in the middle layers.
6 A Theoretical Analysis
We present a theoretical analysis that isolates the key tension in finetuning: adapting to a new task while preserving performance on a base task. We show that full-model finetuning fits the finetuning task perfectly but can catastrophically forget the base task, whereas LoRA—by restricting updates to be low-rank—can simultaneously fit the finetuning task and retain the base task when the finetuning target is low-rank and the base target has large intrinsic dimension. Moreover, the analysis reveals that a higher-dimensional base task is more tolerant to rank overshoot, providing insight into when and why LoRA works in practice.
6.1 Setup
Data and model. We study a family of regression tasks where the goal is to learn a linear map from ${\mathbb{R}}^{d}$ to ${\mathbb{R}}^{k}$. Each task is specified by ${\mathcal{T}}=({\mathcal{V}},s,\sigma,{\bm{\Theta}})$, where ${\mathcal{V}}\subseteq{\mathbb{R}}^{d}$ is a task-specific signal subspace, $s>0$ is the signal scale, $\sigma>0$ is the noise scale, and ${\bm{\Theta}}\in{\mathbb{R}}^{d\times k}$ is the target map. Data are generated by ${\bm{x}}={\bm{z}}+{\bm{\xi}}$ with ${\bm{z}}\sim{\mathcal{N}}(0,s^{2}{\bm{P}}_{{\mathcal{V}}})$ and ${\bm{\xi}}\sim{\mathcal{N}}(0,\sigma^{2}{\bm{P}}_{{\mathcal{V}}^{\perp}})$, and labels are ${\bm{y}}={\bm{z}}^{\top}{\bm{\Theta}}$, where ${\bm{P}}_{{\mathcal{V}}}{\bm{\Theta}}={\bm{\Theta}}$. We consider linear predictors $f_{{\bm{W}}}({\bm{x}})={\bm{x}}^{\top}{\bm{W}}$ trained under the squared loss $L({\bm{W}},{\mathcal{T}})\!=\!\mathbb{E}\bigl[\|{\bm{x}}^{\top}{\bm{W}}\!-\!{\bm{y}}\|^{2}\bigr]$. For comparability across tasks, we fix the label scale such that $\mathbb{E}\|{\bm{y}}\|^{2}=C$ for a fixed constant $C>0$.
Base vs. finetuning tasks. We consider a base task ${\mathcal{T}}_{\rm base}=({\mathcal{V}}_{\rm base},s_{\rm base},\sigma_{\rm base},{\bm{\Theta}}_{\rm base})$ and a finetuning task ${\mathcal{T}}_{\rm ft}=({\mathcal{V}}_{\rm ft},s_{\rm ft},\sigma_{\rm ft},{\bm{\Theta}}_{\rm ft})$, and assume the following:
  1. (A1)
    ${\mathcal{V}}_{\rm base}\perp{\mathcal{V}}_{\rm ft}$, i.e., the signal subspaces of the base and finetuning tasks are orthogonal.
  2. (A2)
    $0<\!\sigma_{\rm base}\!=\!o(s_{\rm ft})$ and $0\!<\!\sigma_{\rm ft}\!=\!o(s_{\rm base})$, i.e., each task has nonzero noise, but its noise is asymptotically much smaller than the other task’s signal.
Additionally, the following spectral separation condition is used only for Theorem 6.3:
  1. (A3)
    $\frac{\sigma_{1}({\bm{\Theta}}_{\rm base})}{\sigma_{r_{\rm ft}}({\bm{\Theta}}_{\rm ft})}<\frac{s_{\rm ft}}{\sigma_{\rm ft}}$, where $\sigma_{i}(\cdot)$ denotes the $i$-th singular value in nonincreasing order. In words, the ratio between the largest singular value of ${\bm{\Theta}}_{\rm base}$ and the smallest nonzero singular value of ${\bm{\Theta}}_{\rm ft}$ is strictly smaller than the signal-to-noise ratio of the finetuning task.
Finetuning from a base model. Finetuning initializes at ${\bm{W}}_{0}={\bm{\Theta}}_{\rm base}$, which is optimal for the base task, and then finetunes the model on the finetuning task. We compare the following approaches. (i) full-model finetuning. We run gradient descent over $\bm{W}$} to minimize $L({\bm{W}},{\mathcal{T}}_{\rm ft})$. Since the population loss is a convex quadratic function of $\bm{W}$} with a unique minimizer, classical results in convex optimization imply that, for a sufficiently small stepsize $0<\eta<\frac{1}{\max\{s_{\mathrm{ft}}^{2},\sigma_{\mathrm{ft}}^{2}\}}$, gradient descent converges linearly to the unique minimizer. We therefore focus on the converged solution, which we denote by
${\bm{W}}_{\mathrm{full}}\in\arg\min_{{\bm{W}}}L({\bm{W}},{\mathcal{T}}_{\mathrm{ft}}),$
and treat it as the final model obtained by full-model finetuning. (ii) LoRA finetuning. We reparameterize the model weights as ${\bm{W}}={\bm{W}}_{0}+{\bm{B}}{\bm{A}}$, where ${\bm{B}}\in{\mathbb{R}}^{d\times r}$ and ${\bm{A}}\in{\mathbb{R}}^{r\times k}$ are trainable parameters, and perform gradient descent over $({\bm{B}},{\bm{A}})$:
$({\bm{B}}_{t+1},{\bm{A}}_{t+1})=({\bm{B}}_{t},{\bm{A}}_{t})-\eta\,\nabla_{{\bm{B}},{\bm{A}}}L({\bm{W}}_{0}+{\bm{B}}_{t}{\bm{A}}_{t},{\mathcal{T}}_{\mathrm{ft}}).$
Under mild conditions on the stepsize and initialization (see detailed discussion in Appendix D.7), gradient descent converges to a global minimizer. We therefore focus on the converged parameters $({\bm{B}}_{\infty},{\bm{A}}_{\infty})$. Equivalently, letting ${\bm{S}}:={\bm{B}}{\bm{A}}$, LoRA restricts the update to satisfy ${\textbf{rank}}{{\bm{S}}}\leq r$. Thus, the final weights obtained by LoRA can be written as
${\bm{W}}_{\mathrm{lora}}={\bm{W}}_{0}+{\bm{S}}^{*},\text{where}\,{\bm{S}}^{*}\in\arg\min_{{\textbf{rank}}{{\bm{S}}}\leq r}L({\bm{W}}_{0}+{\bm{S}},{\mathcal{T}}_{\mathrm{ft}}).$
Key dimensions. We write $r_{\rm ft}:={\textbf{rank}}({\bm{\Theta}}_{\rm ft})$ and define the intrinsic dimension (stable rank)
${\textbf{intdim}}({\bm{M}})\;:=\;\|{\bm{M}}\|_{F}^{2}/\|{\bm{M}}\|_{2}^{2}\;\in[1,{\textbf{rank}}({\bm{M}})].$
Overshoot. We also define the overshoot
$\Delta\;:=\;(r-r_{\rm ft})_{+},$
measuring how much the LoRA budget exceeds the minimal rank needed to represent the finetuning map exactly.
6.2 Comparison and a Deep Dive into LoRA
Figure 8: Numerical experiments under the theoretical setup. We let the singular values of ${\bm{\Theta}}_{\rm base}$ and ${\bm{\Theta}}_{\rm ft}$ follow a power-law decay. We plot both losses against the LoRA rank $r$. ${\textbf{intdim}}({\bm{\Theta}}_{\rm ft})$ is fixed at $1.2$, while ${\textbf{intdim}}({\bm{\Theta}}_{\rm base})$ increases from left to right ($1.2$, $15.5$, and $118.5$). The red shaded region indicates values of $r$ for which both losses are $\leq\!0.05$. As ${\textbf{intdim}}({\bm{\Theta}}_{\rm base})$ grows, this region expands from nonexistent to wide.
Full-model finetuning forgets. Full finetuning achieves the finetuning optimum but overwrites the base solution.
Theorem 6.1 (Full-model finetuning forgets).
Assume (A1)(A2). As proved in Appendix D.3, full finetuning satisfies
$L({\bm{W}}_{\rm full},{\mathcal{T}}_{\rm ft})=0,\qquad L({\bm{W}}_{\rm full},{\mathcal{T}}_{\rm base})\geq C.$
The finetuning task must be low-dimensional for LoRA to work. Even ignoring interference with the base task, a rank-$r$ update can capture at most a rank-$r$ approximation of ${\bm{\Theta}}_{\rm ft}$. The next result shows that the finetuning loss cannot be small unless the intrinsic dimension of ${\bm{\Theta}}_{\rm ft}$ is comparable to the chosen rank $r$. The proof is in Appendix D.4.
Theorem 6.2.
Assume (A1). For any LoRA solution with $r$,
$L({\bm{W}}_{\rm lora},{\mathcal{T}}_{\rm ft})\;\geq\;C\Bigl(1-\frac{r}{{\textbf{intdim}}({\bm{\Theta}}_{\rm ft})}\Bigr)_{+}.$
Consequently, if ${\textbf{intdim}}({\bm{\Theta}}_{\rm ft})\!\gg\!r$, then $L({\bm{W}}_{\rm lora},{\mathcal{T}}_{\rm ft})\gtrsim C$.
Figure 9: Safety finetuning on Qwen2-1.5B-Instruct. Utility is evaluated on instruction-following tasks (COPA, MultiRC, CB), and safety on StrongREJECT. In this non-reasoning, instruction-following setting, LoRA no longer provide an effective safety–utility trade-off.
Interaction between overshoot and the intrinsic dimension of the base task. When $r$ is large enough to represent the finetuning map, LoRA can fit the finetuning task while approximately preserving the base task. However, the degree of base-task degradation depends on the overshoot $\Delta$ and on how spread out the base task is in parameter space. A large ${\textbf{intdim}}({\bm{\Theta}}_{\rm base})$ means the energy of ${\bm{\Theta}}_{\rm base}$ is distributed across many directions, making the base task more tolerant to overshoot. The next result formalizes this: once $r$ is large enough to fit the finetuning map, the finetuning loss becomes negligible, and the base loss scales inversely with ${\textbf{intdim}}({\bm{\Theta}}_{\rm base})$. The proof is in Appendix D.5.
Theorem 6.3 (Near-perfect LoRA finetuning when the base task has large intrinsic dimension).
Assume (A1)(A3). If $r\geq r_{\rm ft}$, then the LoRA-finetuned model satisfies
$L({\bm{W}}_{\rm lora},{\mathcal{T}}_{\rm ft})$ $=o(1),$
$L({\bm{W}}_{\rm lora},{\mathcal{T}}_{\rm base})$ $\leq C\cdot\frac{\Delta}{{\textbf{intdim}}({\bm{\Theta}}_{\rm base})}+o(1).$
Consequently, if ${\textbf{intdim}}({\bm{\Theta}}_{\rm base})=\omega(\Delta)$, then $L({\bm{W}}_{\rm lora},{\mathcal{T}}_{\rm base})=o(1)$ even with overshoot.
Remark 6.4 (Lower bounds and tightness).
The following rank-based lower bound always holds:
$L({\bm{W}}_{\rm lora},{\mathcal{T}}_{\rm base})\;\geq\;C\cdot\frac{\Delta}{{\textbf{rank}}({\bm{\Theta}}_{\rm base})}+o(1).$
Moreover, under a mild flat-spectrum condition (see Appendix D.6), one obtains a matching lower bound of order
$L({\bm{W}}_{\rm lora},{\mathcal{T}}_{\rm base})\;\gtrsim\;C\,\frac{\Delta}{{\textbf{intdim}}({\bm{\Theta}}_{\rm base})}.$
Importantly, Theorem 6.3 and Remark 6.4 show that the sensitivity to overshoot $\Delta$ is governed by ${\textbf{intdim}}({\bm{\Theta}}_{\rm base})$. In practice, it is rarely possible to choose the LoRA rank $r$ perfectly, and since $r$ is discrete, any mismatch yields $\Delta\geq 1$. Therefore, a key requirement for robust performance is that the base task have sufficiently large intrinsic dimension, so that it can tolerate inevitable overshoot.
To illustrate this effect, we conduct numerical experiments that mirror our theoretical setup, setting the singular values of ${\bm{\Theta}}_{\rm base}$ and ${\bm{\Theta}}_{\rm ft}$ to follow a power-law decay. Fig. 8 plots the base and finetuning losses as a function of the LoRA rank $r$ across three regimes, with the intrinsic dimension of the base task increasing from left to right. In the leftmost panel (${\textbf{intdim}}({\bm{\Theta}}_{\rm base})\!=\!1.2$), no choice of $r$ achieves good performance on both tasks. In the middle panel (${\textbf{intdim}}({\bm{\Theta}}_{\rm base})\!=\!15.5$), good performance is achievable, but the base loss is highly sensitive and increases sharply even with a small overshoot; correspondingly, the red “good” region is narrow and requires a very precise choice of $r$. In the rightmost panel (${\textbf{intdim}}({\bm{\Theta}}_{\rm base})\!=\!118.5$), the good region widens substantially, demonstrating increased tolerance to overshoot and making LoRA much less sensitive to the exact choice of $r$.
Taken together: LoRA succeeds with high-rank base tasks and low-rank finetuning tasks. Theorems 6.2 and 6.3 formalize two complementary conditions under which LoRA is effective: the finetuning task must be sufficiently low-rank (otherwise the finetuning loss is large), and the base task must be sufficiently high-rank to tolerate the inevitable overshoot encountered in practice. This perspective helps explain why LoRA is effective in our earlier experiments. In particular, safety finetuning a reasoning model matches this regime: the base capability we aim to preserve—reasoning—is highly complex and thus plausibly “high-rank” in nature. In contrast, the finetuning objective—safety—appears to be relatively low-rank, as suggested by prior work on steering vectors (Panickssery et al., 2023) and “refusal” features (Arditi et al., 2024; Yu et al., 2024), which show that safety-related behavior can be mediated by a single direction or a small number of directions in a model’s internal representations.
To probe this prediction in a more realistic setting, we consider safety finetuning an instruction-tuned (non-reasoning) model, where the base capability is instruction following rather than reasoning. We finetune Qwen2-1.5B-Instruct for safety and evaluate utility on instruction-following tasks (Wang et al., 2019) including COPA, MultiRC and CB, and safety on StrongREJECT. Fig. 9 shows that LoRA is no longer effective in this regime and fails to provide a favorable safety–utility trade-off. This aligns with our theoretical insight that LoRA performs best when the base capability to be preserved is intrinsically high-dimensional. Here, the relevant base capability is instruction following, which is intuitively less complex than reasoning and thus plausibly lower-dimensional, potentially accounting for LoRA’s weaker performance in this setting.